# Puzzle Similarity: A Perceptually-guided Cross-Reference Metric for Artifact Detection in 3D Scene Reconstructions

### v2

@@ 4 Results
(a) Artifact-ridden rendering
@@
Figure 4: Selection of image quality maps for artifact-ridden renderings from various scenes. The last column shows ground-truth human assessments from our collected dataset. Note that our metric provides the finest resolution enabling better artifact segregation.
We will now analyze how our method stacks against competing approaches for both No-Reference and Cross-Reference visual map prediction in the context of reconstruction and image-based rendering. To quantify the correlation between all these maps and human assessment, we present a novel dataset on human artifact identification, which we manually collected and can be found here111www.placeholder__\__fake__\__link__\__for__\__dataset.com to facilitate future research on the topic. As for our method, for each different scene, we compute embeddings on their respective training dataset and leverage each of them respectively to compute visual quality maps on validation and artifact-ridden views, as explained in Sec. 3. CrossScore leverages the same training views for its quality map.
@@ 4.1 A Novel Artifact Identification Dataset
We created a dataset of human-perceived artifacts in 3D reconstructed views with corresponding ground truths collected through a user study. To generate images exhibiting typical reconstruction artifacts, we apply 3D Gaussian Splatting [17] to twelve scenes from the Mip-NeRF360 [4], Tanks and Temples [18], and Deep Blending [10] datasets. We use default parameters but intentionally withhold a significant portion of training images, which we later utilize for additional validation. By omitting these views during training, we increase the likelihood of artifacts appearing in the withheld images. For each dataset, we selected three renderings that demonstrated a mix of well-reconstructed areas, strong artifacts, and subtle artifacts, resulting in 36 samples across 12 datasets.
@@
We evaluate our method against both No-Reference and Cross-Reference IQMs. To assess their alignment with human perception, we correlate their maps with the human ratings from our dataset, as described in Sec. 4.1. NR- and CR-IQMs are the only metrics that can detect artifacts without a direct reference, but the way we collected our dataset gives us access to a ground truth that is normally unavailable. This enables us to assess FR-IQMs and current VDPs too, although they are otherwise not suitable for the objective in question. We include extensive results in the Supplemental and show that, unlike any other competing metric, our metric even outperforms the best FR-IQMs on this benchmark, proving the general superiority of our method.
@@ 4.2 Evaluation
Table 1: Pearson and Spearman Correlation between NR-, and CR-IQMs and Human Perception per Dataset. The dashed line separates NR (above) from CR (below) metrics.
@@
To correlate metric outputs to our human ratings, we first compute each metric map for each sample rendering and then compute the Pearson correlation coefficient (PCC) and Spearman’s rank correlation coefficient (SRCC). To account for the different domains of the compared metrics and possibly non-linear relations, we fit a 5-parameter logistic curve for a fair comparison as suggested by [1, 21, 65, 41]:
@@
where $x$ is an individual quality score and $a_{1\dots 5}$ are optimized through gradient ascent to maximize PCC or SRCC.
@@ Results
Tab. 1 reports the average Pearson and Spearman correlations for each scene. Our dataset includes three images per scene, with varying artifact types per scene. For example, the garden scene has prominent black regions due to holes in the reconstruction, making artifact detection straightforward and leading to high correlation scores for most metrics. However, scenes like treehill, stump, and flowers exhibit artifacts in the form of blurry or unnatural textures while preserving similar color distributions to the ground truth. Puzzle Similarity consistently achieves high correlation with human-perceived artifacts across all datasets, retaining smaller variance across datasets than all comparable methods (See Tab. 2). The higher variance in the averaged results is mainly due to scenes having different types of artifacts, where some are harder to identify than others (e.g., it is easier to identify black areas than small diffuse blobs). Keeping a small variance implies robust performance across various artifact types. We include extended quantitative and qualitative results in our Supplemental.
@@ 4.3 Comparison with No-Reference Metrics
We compare with other No-Reference quality metrics capable of producing spatial quality maps [63, 14, 67, 58, 53]. CNNIQA [14] was applied on patches as described in their paper. We applied padding to avoid cropping the borders and bilinearly upsampled the final map. PIQE [33] already produces three different kinds of maps that we averaged. PAL4VST [67], PaQ-2-PiQ [58] produce maps and were applied as they described in their papers but bilinearly upsampled to match the human maps’ resolution.
@@ 4.4 Comparison with Cross-Reference Metrics
CrossScore is to our knowledge the only other CR metric besides us, also reliant on the training views. We also bilinearly upsampled its output to match human maps. We show comparisons to CrossScore in Tab. 1 and Tab. 2, in Fig. 4 and extended results in the Supplementary. We outperform CrossScore in most datasets and show better performance both on average and in terms of consistency (with a substantially smaller standard deviation among results). While the expensive domain-specific pretraining of CrossScore should, in theory, be superior to our feature-space patch matching leveraging general models pre-trained on collections of natural images, their reliance on SSIM as its target quality assessment metric limits its potential to accurately model human quality assessment, due to the well-known limitations of the metric in this regard [36, 37, 68]. Their DINOv2 encoder limits map resolution to 14x14 blocks, reducing artifact localization accuracy. Furthermore, our approach is markedly simpler, can leverage any CNN as feature encoder, and can be adapted to specific domains by simply changing this backbone, without re-training or distilling any component of the framework.
@@ 4.5 Comparison with Full-Reference Metrics
While unsuitable for the few-shot reconstruction and no-reference artifact identification tasks our work focuses on, we provide extensive comparisons to Full-Reference metrics as well, given that our dataset further provides reference images (which are not normally available for these tasks). On average we further outperform all FR metrics, while CrossScore falls behind FovVideoVDP. Thus, in the context of 3D reconstruction, our metric performs better than all tested NR, CR and FR metrics. Investigating why even the most advanced VDPs, rooted in complex models of low-level human visual processing, quantitatively fall behind us, remains a fascinating avenue for future work.
@@
5 Application: Progressive Automatic Artifact Inpainting
Finally, we showcase an application of our metric in automatic restoration of novel views from a reconstructed scene. Whenever it is possible to establish a visual distribution (e.g. we have a training dataset available), we can recursively use our metric to automatically identify visual outliers in novel views and remove them through inpainting.
@@ 5 Application: Progressive Automatic Artifact Inpainting
We can take a new image $\mathcal{I}$ and employ our PuzzleSim metric to obtain the quality map $\mathcal{Q}$.
@@
To apply neural inpainting, we first need to create a binary mask from the quality map $\mathcal{Q}$, indicating the areas to be inpainted. This involves finding an optimal threshold $\tau$ that clearly distinguishes artifact regions. The effectiveness of inpainting depends on carefully setting this mask. If the mask is too large, the inpainting may inadvertently remove clean parts of the scene. If the mask is too small, artifacts might be left untouched. In order to automatically find a balanced threshold, we use a conservative, iterative approach to refine the test image based on the assumption that artifacts have below-average quality scores. For an initial threshold, we select $N=50$ candidate values, uniformly spaced between the lowest and mean quality scores that we use to threshold the quality map.
$\tau_{i}=$ $\min(\mathcal{Q})+\frac{i}{N-1}(\text{mean}(\mathcal{Q})-\min(\mathcal{Q}))$ (7)
$M^{(h,w)}_{i}=$$\begin{cases}1&\text{if }\mathcal{Q}^{(h,w)}\leq\tau_{i}\\ 0&\text{if }\mathcal{Q}^{(h,w)}>\tau_{i}\end{cases}$
with $i$, $h$, and $w$ representing indices where $i=0,\dots,N-1$, $h=1,\dots,H_{\mathcal{I}}$, and $w=1,\dots,W_{\mathcal{I}}$. With LaMa (big) [44] we generate inpaintings using all $N$ masks and recompute PuzzleSim for each option. The quality of each inpainted candidate is evaluated by calculating the average quality difference within the inpainted region, denoted as $\delta_{i}$. To discourage overly large masks, we add a regularization term that penalizes them. Further details on the mathematical definitions of $\delta_{i}$ and the regularization term are provided in the Supplementary material. We then select the candidate that maximizes $\delta$. After determining the initial threshold, we iteratively refine the inpainted image by drawing new thresholds close to the previous one. For each, we repeat the selection procedure, but with fewer candidates sampled within a narrow range around the previous threshold. The size of this interval depends on a hyperparameter $\alpha$ and the spread of quality scores. Keeping this range small ensures stable convergence and prevents excessive, disrupting inpainting. If the upper limit of the interval is below the minimum quality value $\min_{h,w}\hat{\mathcal{Q}}^{(h,w)}$, we revert to the initial sampling method in Eq. 7 as an empty mask would be meaningless and cause division by zero when computing $\delta_{i}$. Finally, we terminate the process if no further improvement is achieved (i.e., $\max_{i}\delta_{i}\leq 0$), returning the final inpainting result. This framework guarantees a monotonic improvement in PuzzleSim quality. In Fig. 5, we showcase several novel views from the reconstructed scenes garden, kitchen, and flowers using only a fraction of the original training views (20-30%). We process this artifact-ridden new view through the iterative inpainting framework presented above. Our method successfully detects and inpaints artifacts in the original reconstruction, producing high-quality inpainting consistent with the distribution of the original scenes.
@@ 6 Limitations and Future Work
While our method demonstrates promising results, there are some limitations to consider. Even with our optimized implementation, finding the maximum similarity for a great number of vectors becomes expensive as the number of reference images and image resolution rises (see Supplemental material for implementation and runtime analysis). Performing approximate maximum search or fitting Gaussian mixture models in the embedding space can improve computational performance [66, 13]. Furthermore, our metric is empirically calibrated, but choosing the weights to combine layers and weighting in the channel dimension in a data-driven manner could advance the metric further. The resolution at which our metric can be utilized is currently bound by the CNN backbone’s generalizability to higher resolutions. Currently, we can not support Vision Transformer backbones as it takes special care to maintain a limited receptive field. Although our metric is differentiable, it is unlikely to produce valuable gradients due to the max operation across many vectors. The next step would be to explore softmax alternatives to make the metric more suitable for gradient-based optimization.
@@ 7 Conclusion
We proposed Puzzle Similarity, a Cross-Reference visual quality metric for detecting and localizing artifacts in novel views of 3D scene reconstructions. By leveraging learned patch statistics from input views, our method generates spatial artifact maps without requiring ground-truth references, addressing a key challenge in evaluating reconstructed scenes. To enable the evaluation of Cross-Reference metrics, we also provide a dataset of human-assessed quality and artifact localization for 3D scene reconstruction.
Our evaluation shows that Puzzle Similarity outperforms all tested Full-Reference, Cross-Reference and No-Reference metrics in capturing artifacts aligned with human perception, demonstrating robustness across diverse artifact types and texture-rich scenes without requiring direct references. Furthermore, we apply our metric to automatic image restoration, illustrating its potential to enhance scene reconstruction quality. Puzzle Similarity provides an effective, perceptually aligned, reference-free solution for artifact localization, with promising applications in few-shot reconstruction and guided acquisition.

### v3

@@ 4 Results
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Artifact-ridden view
@@
Figure 4: Selection of image quality maps for artifact-ridden renderings from various scenes. The last column shows ground-truth human assessments from our collected dataset. Note that our metric provides the finest resolution, enabling better artifact segregation.
We will now analyze how our method compares against competing approaches for both no-reference and cross-reference visual map prediction in the context of reconstruction and image-based rendering. To quantify the correlation between all these maps and human assessment, we present a novel dataset on human artifact identification, which we manually collected and can be found here [^1] https://huggingface.co/datasets/nihermann/annotated-3DGS-artifacts to facilitate future research on the topic. As for our method, for each different scene, we compute embeddings on their respective training dataset to compute similarity maps on test views, as explained in Sec. 3. CrossScore leverages the same set of training views for its map predictions.
@@ 4.1 A Novel Artifact Identification Dataset
We created a dataset of human-perceived artifacts in 3D reconstructed views with corresponding ground truths collected through a user study. To generate images exhibiting typical reconstruction artifacts, we apply 3D Gaussian Splatting [15] to twelve scenes from the Mip-NeRF360 [3], Tanks and Temples [16], and Deep Blending [10] datasets. We use default parameters but intentionally withhold a significant portion of training images, which we later utilize for additional validation. By omitting these views during training, we increase the likelihood of artifacts appearing in the withheld views. For each dataset, we selected three renderings that demonstrated a mix of well-reconstructed areas, strong artifacts, and subtle artifacts, resulting in 36 samples across 12 datasets.
@@
We evaluate our method against both no-reference and cross-reference metrics. To assess their alignment with human perception, we correlate their maps with the human ratings from our dataset, as described in Sec. 4.1. NR and CR metrics are the only metrics that can detect artifacts without a direct reference, but the way we collected our dataset gives us access to a ground truth that is normally unavailable. This enables us to assess FR metrics and current VDPs too, although they are otherwise not suitable for the objective in question. We include extensive results in the Supplemental and show that, unlike any other competing metric, our metric even outperforms the best FR metrics on this benchmark, proving the general superiority of our method.
@@ 4.2 Evaluation
Table 1: Pearson and Spearman Correlation between NR, and CR metrics and Human Perception per Dataset. The dashed line separates NR (above) from CR (below) metrics.
@@
To correlate metric outputs to our human segmentations, we first compute each metric map for each rendering and then compute the Pearson correlation coefficient (PCC) and Spearman’s rank correlation coefficient (SRCC). To account for the different domains of the compared metrics and possibly non-linear relations, we fit a 5-parameter logistic curve for a fair comparison as suggested by [1, 19, 60, 34]:
@@
where $x$ is an individual pixel score and $a_{1\dots 5}$ are optimized through gradient ascent to maximize PCC or SRCC.
@@ Results
Tab. 1 reports the average Pearson and Spearman correlations for each scene. Our dataset includes three images per scene, with varying artifact types per scene. For example, the garden scene has prominent black regions due to holes in the reconstruction, making artifact detection straightforward and leading to high correlation scores for most metrics. However, scenes like treehill, stump, and flowers exhibit artifacts in the form of blurry or unnatural textures while preserving similar color distributions to the ground truth. Puzzle Similarity consistently achieves high correlation with human-perceived artifacts across all datasets, retaining smaller variance across datasets than all competitive methods (See Tab. 2). The higher variance in the averaged results is mainly due to scenes having different types of artifacts, where some are harder to identify than others (e.g., it is easier to identify black areas than small diffuse blobs). Keeping a small variance implies robust performance across various artifact types. We include extended quantitative and qualitative results in our Supplemental.
@@ 4.3 Comparison with No-Reference Metrics
We compare with other no-reference metrics capable of producing spatial visibility maps [58, 13, 61, 53, 47]. CNNIQA [13] was applied on patches as described in their paper. We applied padding to avoid cropping the borders and bilinearly upsampled the final map. PIQE [29] already produces three different kinds of maps that we averaged. PAL4VST [61], PaQ-2-PiQ [53] produce maps and were applied as described in their papers, but bilinearly upsampled to match the human maps’ resolution.
@@ 4.4 Comparison with Cross-Reference Metrics
CrossScore is, to our knowledge, the only other CR metric besides ours that is also reliant on the training views. We also bilinearly upsampled its output to match the human maps’ resolution. We show comparisons to CrossScore in Tab. 1 and Tab. 2, in Fig. 4 and extended results in the Supplementary. We outperform CrossScore on most datasets and show better performance both on average and in terms of consistency (with a substantially smaller standard deviation among results). While the expensive domain-specific pretraining of CrossScore should, in theory, be superior to our feature-space patch matching leveraging general models pre-trained on collections of natural images, their reliance on SSIM as its target quality assessment metric limits its potential to accurately model human quality assessment, due to the well-known limitations of the metric in this regard [31, 32, 62]. Their DINOv2 encoder limits map resolution to 14x14 blocks, reducing artifact localization fidelity. Furthermore, our approach is notably simpler: we can leverage any CNN as a feature encoder, allowing seamless adaptation to specific domains simply by swapping out the backbone. No retraining or distillation of any framework component is required.
@@ 4.5 Comparison with Full-Reference Metrics
Although full-reference metrics, unlike our method, require a direct reference image for detecting artifacts, we also provide an extensive comparison to them based on our dataset, which includes reference images. On average, we outperform all FR metrics, while CrossScore falls behind FovVideoVDP. Thus, in the context of 3D reconstruction, our metric performs better than all tested NR, CR, and FR metrics. Investigating why even the most advanced VDPs, rooted in complex models of low-level human visual processing, quantitatively fall behind remains a fascinating avenue for future work.
@@
5 Application: Progressive Inpainting
Finally, we showcase an application of our metric in automatic restoration of novel views from a reconstructed scene. Whenever it is possible to establish a visual distribution (e.g., we have a training dataset available), we can recursively use our metric to automatically identify visual outliers in novel views and remove them through inpainting. In the Supplementary, we present a quantitative ablation study to show that our metric performs best in this application.
@@ 5 Application: Progressive Inpainting
We can take a new image $\mathcal{I}$ and employ our PuzzleSim metric to obtain the similarity map $\mathcal{S}$.
@@
To apply neural inpainting, we first need to create a binary mask from the similarity map $\mathcal{S}$, indicating the areas to be inpainted. This involves finding an optimal threshold $\tau$ that clearly distinguishes outlier regions. The effectiveness of inpainting depends on carefully setting this mask. If the mask is too large, the inpainting may inadvertently remove clean parts of the scene. If the mask is too small, artifacts might be left untouched. In order to automatically find a balanced threshold, we use a conservative, iterative approach to refine the test image based on the assumption that artifacts have below-average similarity scores. For an initial threshold, we select $N=50$ candidate values, uniformly spaced between the lowest and mean similarity scores that we use to threshold the similarity map.
$\tau_{i}=$ $\min(\mathcal{S})+\frac{i}{N-1}(\text{mean}(\mathcal{S})-\min(\mathcal{S}))$ (7)
[…]
with $i$, $h$, and $w$ representing indices where $i=0,\dots,N-1$, $h=1,\dots,H_{\mathcal{I}}$, and $w=1,\dots,W_{\mathcal{I}}$. With LaMa (big) [37] we generate inpaintings using all $N$ masks and recompute PuzzleSim for each option. The quality of each inpainted candidate is evaluated by calculating the average similarity difference before and after inpainting, denoted as $\delta_{i}$. To discourage overly large masks, we add a regularization term that penalizes them. Further details on the mathematical definitions of $\delta_{i}$ and the regularization term are provided in the Supplementary material. We then select the candidate that maximizes $\delta$. After determining the initial threshold, we iteratively refine the inpainted image by sampling new thresholds close to the previous one. The size of this interval depends on a hyperparameter $\alpha$ and the spread of similarity scores. Keeping this range small ensures stable convergence and prevents excessive, disrupting inpainting. If the upper limit of the interval is below the minimum similiarity value $\min_{h,w}\hat{\mathcal{S}}^{(h,w)}$, we revert to the initial sampling method in Eq. 7 as an empty mask would be meaningless and cause division by zero when computing $\delta_{i}$. Finally, we terminate the process if no further improvement is achieved (i.e., $\max_{i}\delta_{i}\leq 0$), returning the final inpainting result. This framework guarantees a monotonic improvement in PuzzleSim similarity. In Fig. 5, we showcase several novel views from the reconstructed scenes garden, kitchen, and flowers using only a fraction of the original training views (20-30%). We process this artifact-ridden new view through the iterative inpainting framework presented above. Our method successfully detects and inpaints artifacts in the original reconstruction, producing high-quality inpainting consistent with the distribution of the original scenes.
@@ 6 Limitations and Future Work
While our method demonstrates promising results, there are some limitations to consider. Even with our optimized implementation, finding the maximum similarity for a great number of vectors becomes expensive as the number of reference images and image resolution rise (see Supplemental material for implementation and runtime analysis). Performing approximate maximum search or fitting Gaussian mixture models in the embedding space can improve computational performance [60, 12]. Furthermore, our metric is empirically calibrated, but choosing the weights to combine layers and weighting in the channel dimension in a data-driven manner could advance the metric further. The resolution at which our metric can be utilized is currently bound by the CNN backbone’s generalizability to higher resolutions. Currently, we can not support Vision Transformer backbones as it takes special care to maintain a limited receptive field. Although our metric is differentiable, it is unlikely to produce valuable gradients due to the max operation across many vectors. Alternatively, softmax operations could be explored to make the metric more suitable for gradient-based optimization.
@@ 7 Conclusion
We proposed Puzzle Similarity, a cross-reference image metric for detecting and localizing artifacts in novel views of 3D scene reconstructions. By leveraging learned patch statistics from input views, our method generates spatial artifact maps without requiring ground-truth references, addressing a key challenge in evaluating reconstructed scenes. To enable the evaluation of cross-reference metrics, we also provide a dataset of human-assessed quality and artifact localization for 3D scene reconstruction.
Our evaluation shows that Puzzle Similarity outperforms all tested full-reference, cross-reference and no-reference metrics in capturing artifacts aligned with human perception, demonstrating robustness across diverse artifact types and texture-rich scenes. Furthermore, we apply our metric to automatic image restoration, illustrating its potential to enhance scene reconstruction quality. Puzzle Similarity provides an effective, perceptually aligned, reference-free solution for artifact localization, with promising applications in few-shot reconstruction and guided acquisition.